OpenStudy (quinton1012):

Shalika bought a purse for $120. The tax rate is 9%. Which proportion can be used to find the amount of tax Shalika will pay? i will medal

OpenStudy (quinton1012):

@misty1212

OpenStudy (quinton1012):

@YumYum247

OpenStudy (quinton1012):

@Error1603

OpenStudy (shaik0124):

9% id the tax of $120 so find out how much is 9% of 120

OpenStudy (shaik0124):

tax amount =(120) *9/100

OpenStudy (shaik0124):

10.8 is the tax amount

OpenStudy (quinton1012):

thx

OpenStudy (shaik0124):

did u find ur answer

OpenStudy (quinton1012):

i think

OpenStudy (shaik0124):

what are the options

OpenStudy (quinton1012):

A 9/120=x/100 B 120/x=100/9 C x/120=100/9 D 120/9=100/x

OpenStudy (quinton1012):

@shaik0124

OpenStudy (quinton1012):

i'm thinking a or d

OpenStudy (shaik0124):

b

OpenStudy (quinton1012):

oh okay thank you!! :D